Two bodies of masses 100 kg and 1000 kg are at a distance 1 m apart.Calculate the gravitational field intensity and potential at the middle point of the line joining them

Grade:11

1 Answers

nive
5 years ago
et at x distance from m₁ and 1 -x distance from m₂ Gravitational field intensity will be zero. 
we know, 
Magnitude Gravitational field intensity = GM/R² 
Here, 
m₁ = 10 Kg , m₂ = 1000 Kg 
A/C to question, 
Gm₁/x² = Gm₂/(1 - x)² 
10/x² =1000/(1 - x)² 
1/x²= 100/(1 - x)² 
1/x = 10/(1 - x) 
1 - x = 10x 
1 = 11x ⇒ x = 1/11 m 
Hence, from m₁ distance will be 1/11 m 
And from m₂ , distance will be 10/11 m